# Think Silently, Think Fast: Dynamic Latent Compression of LLM Reasoning Chains

### v4

@@ 4.2 Comparison to baseline methods on GSM datasets
Furthermore, CoLaR exhibits robust out-of-domain generalization capabilities compared to other latent-based baselines, particularly on the MultiArith dataset. On this dataset, CoLaR experiences minimal performance degradation compared to CoT, while other latent-based methods suffer significantly larger drops. We also provide a detailed case study in Appendix B.
4.3 Ablation studies of CoLaR on GSM datasets
@@ 4.4 Reinforcement learning results on the MATH dataset
We train and evaluate CoLaR with RL on the challenging MATH dataset, using two base models Guo et al. (2025); Yang et al. (2024); Grattafiori et al. (2024). The results are presented in Table 2. Our analysis of the results yields three key conclusions:
@@
(3) Base model quality impacts RL effectiveness. Supervised fine-tuning on CoT resulted in varying performance across the two base LLMs. Meanwhile, CoLaR also demonstrates a significantly larger performance gain on RL when using the higher-quality Qwen-1.5B compared to Llama-1B. This observation aligns with the findings of Gandhi et al. (2025), which suggests that RL substantially activates inherent reasoning capabilities, indicating the importance of base model quality.
We also observed that during the RL training process, CoLaR tends to think longer initially with a rapid rise in accuracy, followed by a phase of thinking shorter accompanied by a more stable increase in accuracy, aligning with the discussion in Section 3.3. Due to space constraints, the detailed training curves are provided in Appendix D.
@@ 5 Limitations
While CoLaR demonstrates superior effectiveness and efficiency in latent reasoning, we acknowledge two important limitations: (1) Like most latent-based methods, the overall performance of CoLaR currently approximates explicit CoT reasoning without surpassing it. We attribute this primarily to our compression-targeted training approach, which focuses on imitating the behavior of explicit CoT. However, we believe that more sophisticated reinforcement learning algorithms that encourage models to explore and exploit diverse reasoning pathways could potentially achieve performance that exceeds traditional CoT methods. (2) We observe that CoLaR struggles to generalize to non-integer compression factors (e.g., $c=1.5$) or to values greater than the maximum training compression factor $r_{max}$. This limitation stems primarily from the discrete tokenization constraints inherent to large language models, which restrict the continuous representation of compression factors. (3) We do not observe significant performance gain applying RL on GSM datasets. We mainly attribute this to that these simple math questions requires less exploration. (4) Beyond technical limitations, our work on enhancing reasoning capabilities in LLMs has significant societal implications. On the positive side, CoLaR could significantly boost the efficiency of existing LLM services. However, potential negative impacts include the risk of amplifying existing biases in reasoning processes and possible misuse for generating more convincing misinformation. To mitigate these risks, we recommend careful monitoring of downstream applications.
@@ Appendix A More implementation details
Computational resources. For the SFT stage experiments, we use a single Ubuntu machine equipped with eight NVIDIA A100 GPUs, implementing automatic mixed-precision in BF16. CoLaR converges on the GSM8k-Aug training set after approximately 40k training steps, requiring about 6 hours of computation. For the RL stage, we use a single A100 GPU to train CoLaR in FP32 precision, with convergence occurring after approximately 4k steps and 12 hours of training.
Figure 5: A case study on the GSM8k validation set. We set the compression factor $c$ to 2 and 5, which produce two latent reasoning chains in length 7 and 3, respectively. We then retrieve tokens with the predicted latents by embedding cosine similarity, and underscore those informative tokens.
Appendix B Case study on latent retrieval
We conduct a case study of our trained CoLaR on the GSM8k validation set. As illustrated in Figure 5, the latents predicted by CoLaR successfully capture dense semantics. Notably, less informative tokens such as “<<” are ignored when applying a higher compression factor $c$, demonstrating both the effectiveness and efficiency of our method.
@@ Appendix B Case study on latent retrieval
Appendix C Layer-wise analyses on compression factors
We further investigate how the compression factor $c$ influences activation patterns across LLM layers, with results shown in Figure 6. Specifically, we tested CoLaR on the same sample as in Section B with compression factors ranging from 1 to 5, calculating the relative activation norm differences between consecutive LLM layers.
@@ Appendix C Layer-wise analyses on compression factors
In contrast, higher compression factors enable CoLaR to process information more densely, with each latent representation carrying richer semantic content. This requires deeper layers to remain actively engaged in analyzing the condensed information and predicting subsequent compressed latents, thereby making more efficient use of the model’s computational capacity. These findings align with observations from previous work on internal thinking processes in transformer models Chen et al. [2025], Saunshi et al. [2025].
@@
Appendix D RL training curves
@@
Appendix E Scaling properties of CoLaR

### v5

@@ 4.2 Comparison to baseline methods on GSM datasets
Furthermore, CoLaR exhibits robust out-of-domain generalization capabilities compared to other latent-based baselines, particularly on the MultiArith dataset, where CoLaR shows minimal performance degradation compared to CoT, while other latent-based methods suffer significant drops.
4.3 Ablation studies of on GSM datasets
@@
Figure 3: A case study on the GSM-8k validation set. We set the compression factor $c$ to 2 and 5, which produce two latent reasoning chains in length 7 and 3, respectively. We then retrieve tokens with the predicted latents by embedding cosine similarity, and underscore those informative tokens.
@@ 4.4 Reinforcement learning results on the MATH dataset
We train and evaluate CoLaR with RL on the challenging MATH dataset, using two base models r1 ; qwen25 ; llama3 . The results are presented in Table 2. Our analysis of the results yields three key conclusions:
@@
(3) Base model quality impacts RL effectiveness. Supervised fine-tuning on CoT resulted in varying performance across the two base LLMs. Meanwhile, CoLaR also demonstrates a significantly larger performance gain on RL when using the higher-quality Qwen-1.5B compared to Llama-1B. This observation aligns with the findings of gandhi2025cognitive , which suggests that RL substantially activates inherent reasoning capabilities, indicating the importance of base model quality.
We also observed that during the RL training process, CoLaR tends to think longer initially with a rapid rise in accuracy, followed by a phase of thinking shorter accompanied by a more stable increase in accuracy, aligning with the discussion in Section 3.3. Due to space constraints, the detailed training curves are provided in Appendix C.
@@
4.5 Scaling CoLaR on larger base model
We scale CoLaR to the 1-Billion and 8-Billion parameter Llama checkpoints to verify that its compression and accuracy gains grow with model capacity. The 8B model also matches the backbone used by TokenSkip, allowing a direct comparison under its published protocol. Experiments are run on GSM-8k, MATH-500 which are slightly different to Section 4.1 to align with TokenSkip’s settings) and the out-of-domain GPQA benchmark (chemistry, biology and physics). The results are shown in Table 3. Moving from 1 B to 8 B parameters raises CoLaR’s accuracy on every dataset while keeping the reasoning chain roughly half the length of the CoT teacher. Reinforcement learning delivers further improvements: +1.1% on GSM-8k, +6.6% on MATH-500 and +5.1% on GPQA, again with large additional compression. Most notably, CoLaR-8B-RL reaches 37.5% on GPQA, surpassing its 35.7% CoT teacher and shortening the reasoning chain by 69%, confirming that latent RL remains effective as models grow.
Table 3: Experimental results of CoLaR-2 and TokenSkip on GSM-8k, MATH-500, and GPQA.
Acc. (%)/#L CoT-1B CoLaR-1B CoT-8B CoLaR-8B CoLaR-8B-RL TokenSkip-8B
GSM-8k 47.5/92.6 40.4/49.2 76.5/93.6 70.8/47.3 71.9/13.3 78.2/113
MATH-500 28.6/176 19.8/84.6 54.6/168 45.8/67.7 52.4/17.6 40.2/292
GPQA 26.4/232 26.4/84.1 35.7/216 32.4/101 37.5/66.7 -
4.6 Interpreting chains of latent thoughts
To quantatively illustate the effectiveness of our proposed CoLaR, and make the latent reasoning process more transparent, we conduct a case study of CoLaR on the GSM8k validation set. The results are illlustrated in Figure 3.
As latent variables are fundamentally scaled sum of word token embeddings, we could directly calculate the cosine simility between latent variables and the enitre LLM embedding matrix:
  • When prompted with compression factor $c=2$, CoLaR auto-regressively produces seven latent variables, and then automatically terminates reasoning process. By calculating similarity scores, each latent variable is capable of retrieving meaningful words such as “<<”, “21”, and the entire chain of latent thoughts could be interpretered as “<<21/7=3> <<5*3=15>”, which exactly matches the correct calculation process.
  • When prompted with higher compression factor $c=5$, less informative tokens such as “<<” are ignored, demonstrating both the effectiveness and efficiency of our proposed dynamic compression mechanism.
@@ 5 Limitations
While CoLaR demonstrates superior effectiveness and efficiency in latent reasoning, we acknowledge two important limitations: (1) Though CoLaR surpassing its CoT teacher on GPQA, on the remaining benchmarks, the overall performance of CoLaR currently approximates explicit CoT reasoning without surpassing it. (2) We observe that CoLaR struggles to generalize to non-integer compression factors (e.g., $c=1.5$) or to values greater than the maximum training compression factor $r_{max}$. This limitation stems primarily from the discrete tokenization constraints inherent to large language models, which restrict the continuous representation of compression factors. (3) Beyond technical limitations, our work on enhancing reasoning capabilities in LLMs has significant societal implications. On the positive side, CoLaR could significantly boost the efficiency of existing LLM services. However, potential negative impacts include the risk of amplifying existing biases in reasoning processes and possible misuse for generating more convincing misinformation. To mitigate these risks, we recommend careful monitoring of downstream applications.
@@ 6 Conclusion
Acknowledgements. This work is supported by the Beijing Natural Science Foundation (L233008) and Xiaomi Inc. We acknowledge the anonymous reviewers for their helpful comments.
@@ Appendix A More implementation details
Appendix B Layer-wise analyses on compression factors
We further investigate how the compression factor $c$ influences activation patterns across LLM layers, with results shown in Figure 6. Specifically, we tested CoLaR on the same sample as in Section 4.6 with compression factors ranging from 1 to 5, calculating the relative activation norm differences between consecutive LLM layers.
@@ Appendix B Layer-wise analyses on compression factors
In contrast, higher compression factors enable CoLaR to process information more densely, with each latent representation carrying richer semantic content. This requires deeper layers to remain actively engaged in analyzing the condensed information and predicting subsequent compressed latents, thereby making more efficient use of the model’s computational capacity. These findings align with observations from previous work on internal thinking processes in transformer models [2, 21].
@@
Appendix C RL training curves
@@
Appendix D Scaling properties of CoLaR